OpenStudy (anonymous):

Solve for x. x^2 +18+4=0

OpenStudy (anonymous):

\[x^2+18x=-4\] \[(x+9)^2=-4+81\] \[(x+9)^2=77\] \[x+9=\pm\sqrt{77}\] \[x=-9\pm\sqrt{77}\]

OpenStudy (anonymous):

completing the square here easier than quadratic formula because the "middle term" 18x has an even coefficient, so you do not have to simplify at the end

OpenStudy (anonymous):

how do I make the plus minus sign with the keyboard?

OpenStudy (anonymous):

\pm

OpenStudy (anonymous):

i mean that is how i do it here